OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (−2, −1) and has a slope of 5. y − 1 = 5(x − 2) y + 1 = 5(x + 2) y + 2 = 5(x + 1) y − 2 = 5(x − 1)

OpenStudy (anonymous):

will give medal

OpenStudy (igreen):

Use the point-slope formula. \(\sf y - y_1 = m(x - x_1)\) Where \(\sf m\) is the slope and \(\sf (x_1, y_1)\) is a point on the line. Can you plug in the given information into the formula?

OpenStudy (anonymous):

y-1=5(x-2) ?

OpenStudy (igreen):

Not exactly. Remember, we have (-2, -1), not (2, 1) \(\sf y - (-1) = 5(x - (-2)\) Can you simplify that?

OpenStudy (anonymous):

OHH

OpenStudy (anonymous):

It turns into positive instead of negative?

OpenStudy (igreen):

Correct!

OpenStudy (igreen):

So what do we have?

OpenStudy (anonymous):

YAY! Hehe thanks :)

OpenStudy (igreen):

No problem \(\huge\ddot\smile\)